I saw this problem in my textbook:
A rectangular pool is surrounded by a walk 4 meters wide. The pool is 6 meters longer than its width. If the total area is 576 square meters more than the area of the pool, find the dimensions of the pool.
How do they get the answer? I've tried by making Total Area = x^2+6x+576, but have been unable to obtain the answer.

Answer by scott8148(6628)

"The pool is 6 meters longer than its width" ___ L = w+6

for the pool ___ A = Lw = w^2 + 6w

"A rectangular pool is surrounded by a walk 4 meters wide"
___ so the width of the total area is w+8 (4 on each side) , and the length is w+14

w^2 + 6w + 576 = w^2 + 22w + 112

464 = 16w ___ 29 = w ___ L = 35